OpenStudy (jozelynw):

I need help............ I don't know how to convert f(n)=f(n-1)+d into the sequence notation a(subscript) n=a+d(n-1)

OpenStudy (jozelynw):

@dan815

jimthompson5910 (jim_thompson5910):

is that all it says? or is there more that's given?

OpenStudy (jozelynw):

No it says "Use your scenario to write the function for the 7th term in your sequence using sequence notation".

OpenStudy (jozelynw):

this is what i filled in Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for 20 minutes on Monday. Every day she increases her practice time by 4 . If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation.

OpenStudy (jozelynw):

20 mins , increases , time by 4 that's what i added

jimthompson5910 (jim_thompson5910):

so they want you to make up a sequence of your own? or can we use this example?

OpenStudy (jozelynw):

they wanted me to make up my own based upon what i filled in and the formula above with the a

jimthompson5910 (jim_thompson5910):

ok what sequence did you make up on your own?

OpenStudy (jozelynw):

a(subscript) n=a+d(n-1)

OpenStudy (jozelynw):

that is what help on filling in on this equation. In the f(n) version it made up one which is f(7)=f(6-1)+4

OpenStudy (jozelynw):

when i fill in the sequence equation it looks like this a(subscript)7=a+4(7-1)

jimthompson5910 (jim_thompson5910):

it says `Jasmine practices the piano for 20 minutes on Monday` so the first term of this sequence is 20 making a = 20

jimthompson5910 (jim_thompson5910):

the next day, tuesday, would be a+d = 20+4 = 24 min ------------------------- wednesday would be 24+4 = 28 etc etc

OpenStudy (jozelynw):

so how would that be written

jimthompson5910 (jim_thompson5910):

a = 20 and d = 4 they are plugged into `a+d(n-1)` to get the general nth term

jimthompson5910 (jim_thompson5910):

so, a+d(n-1) = 20+4(n-1) = ???

OpenStudy (jozelynw):

thank you for you help.

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (jozelynw):

@jim_thompson5910 need you help again with another 1

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (jozelynw):

Use your scenario to write the formula for the 5th term in your sequence using sequence notation.

OpenStudy (jozelynw):

the equation for geometric sequence is an=a1•rn−1.

OpenStudy (jozelynw):

my info is Anthony goes to the gym for 45 minutes on Monday. Every day he increases his gym time by 5 minutes.

jimthompson5910 (jim_thompson5910):

`Anthony goes to the gym for 45 minutes on Monday. Every day he increases his gym time by 5 minutes.` this is an ARITHMETIC sequence since the amount is increasing by a fixed amount. d = 5 is the amount going up each time

jimthompson5910 (jim_thompson5910):

if you said maybe something like "increases by 5%", then it would work

OpenStudy (jozelynw):

i see what you're saying , I knew something wasn't right. So change it to 5%

OpenStudy (jozelynw):

then what would i do

jimthompson5910 (jim_thompson5910):

the first term is a = 45 the common ratio is r = 0.05 (decimal form of 5%)

jimthompson5910 (jim_thompson5910):

no sorry, the common ratio would be 1.05 1+0.05 = 1.05

OpenStudy (jozelynw):

so it would be a5=45 x1.05(7-1)

jimthompson5910 (jim_thompson5910):

first term: 45 second term = (first term)*(common ratio) second term = (45)*(1.05) second term = 47.25 etc etc

OpenStudy (jozelynw):

change the 5% to 1.05%?

jimthompson5910 (jim_thompson5910):

I added on 1 because increasing by 5% is the same as multiplying by 1.05

jimthompson5910 (jim_thompson5910):

the nth term should be \[\Large a(r)^{n-1}\] then you plug in a = 45 and r = 1.05

OpenStudy (jozelynw):

ok

OpenStudy (jozelynw):

i have to use that exact equation above \[a _{4=45x 1.05(4}\]

OpenStudy (jozelynw):

the x= times

jimthompson5910 (jim_thompson5910):

the exponent should be n-1 = 4-1 so \[\LARGE a_4 = 45*(1.05)^{4-1}\]

OpenStudy (jozelynw):

well no, because i want to find the 5th term. I simplified it so I didn't put the -1 i just did it.

OpenStudy (jozelynw):

so 5-1

jimthompson5910 (jim_thompson5910):

oh, then you want to compute \(\LARGE a_5\) and not \(\LARGE a_4\)

OpenStudy (jozelynw):

yes it's a5 not a4

OpenStudy (jozelynw):

they wanted me to find a5 the fifth term.

jimthompson5910 (jim_thompson5910):

I see

OpenStudy (jozelynw):

ok so the answer would be what i said rightttt

jimthompson5910 (jim_thompson5910):

yes, it would be \[\LARGE a_5 = 45*(1.05)^{4}\]

OpenStudy (jozelynw):

Thank you, I may need help in like 3 mins again 1 more

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (jozelynw):

a_(5=45 ×〖1.05〗^(5-1) ) @jim_thompson5910

OpenStudy (jozelynw):

Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

jimthompson5910 (jim_thompson5910):

I think you meant to write \[\LARGE a_5 = 45 \times 1.05^{5-1}\] right?

OpenStudy (jozelynw):

Yes, but i don't know how to make it look like that, how do you do that?

OpenStudy (jozelynw):

the geometric series formula Sn= a1−a1rn 1−r

jimthompson5910 (jim_thompson5910):

you can use the equation editor (click the equation button below the text box) I typed in `\LARGE a_5 = 45 \times 1.05^{5-1}` surrounded by `\[` and `\]`

jimthompson5910 (jim_thompson5910):

as for the summation formula, it would be \[\Large S_n = a*\frac{1-r^n}{1-r}\]

jimthompson5910 (jim_thompson5910):

which is equivalent to the form you wrote

OpenStudy (jozelynw):

where the a is by itself it should be with the 1 at the top

jimthompson5910 (jim_thompson5910):

yes the two are equivalent \[\Large S_n = a*\frac{1-r^n}{1-r} = \frac{a-a*r^n}{1-r}\]

OpenStudy (jozelynw):

I'm going to screenshot it

jimthompson5910 (jim_thompson5910):

yeah which is what I wrote here \[\Large S_n = \frac{a-a*r^n}{1-r}\]

jimthompson5910 (jim_thompson5910):

there are 2 versions of the same thing really

OpenStudy (jozelynw):

ok

jimthompson5910 (jim_thompson5910):

they want you to add up a bunch of terms and use the Sn formula to do it quickly

jimthompson5910 (jim_thompson5910):

so for instance, add up the first 8 terms and compute S8

OpenStudy (jozelynw):

So I"m using what you are saying as a example, because Idon't want to seem as a cheater

jimthompson5910 (jim_thompson5910):

you can use another value of n

OpenStudy (jozelynw):

ok 5

OpenStudy (jozelynw):

Oh it says use the same scenario

OpenStudy (jozelynw):

from part 2

jimthompson5910 (jim_thompson5910):

so a = 45 and r = 1.05

OpenStudy (jozelynw):

right ,but i need to make a question now

OpenStudy (jozelynw):

Maybe I can say Anthony goes to the gym at 10 am for 45mins, then IDK what im doing

jimthompson5910 (jim_thompson5910):

so you have to make a completely different scenario?

OpenStudy (jozelynw):

no just use the scenario from part 2 to make up a question

OpenStudy (jozelynw):

i just seen an example on my math lesson so here is something similar "find the sum of the first 5 terms of the geometric series

OpenStudy (jozelynw):

but I don't know what numbers are suppose to be in the geometric series here is the example

jimthompson5910 (jim_thompson5910):

ok so you'll plug a = 45 and r = 1.05 into that Sn formula then you pick any n value you want (you picked n = 5 I think) and plug that in as well

jimthompson5910 (jim_thompson5910):

\[\Large S_n = \frac{a-a*r^n}{1-r}\] \[\Large S_5 = \frac{45-45*1.05^5}{1-1.05}\] \[\Large S_5 = ???\]

OpenStudy (jozelynw):

is it like the example?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (jozelynw):

ok than so what would be the geometric series than for the question?

jimthompson5910 (jim_thompson5910):

I'm not sure what you're asking exactly?

OpenStudy (jozelynw):

because i need to write a question; in the example question in the lesson it has something in the question that says geometric series look at the screenshot again

OpenStudy (jozelynw):

I think it means 3 numbers that have a common ratio of 5%

jimthompson5910 (jim_thompson5910):

you mean they want you to add up the first three terms of the geometric sequence?

OpenStudy (jozelynw):

idk honestly. I'm just listening to you

OpenStudy (jozelynw):

does it look like what it mean in the picture

OpenStudy (jozelynw):

this is a project and it doesn't really explain much

jimthompson5910 (jim_thompson5910):

yeah I wish it had clearer instructions. I think they want you to add up a bunch of terms so say you wanted to add up the first 7 terms a = 45 r = 1.05 n = 7 \[\Large S_n = \frac{a-a*r^n}{1-r}\] \[\Large S_7 = \frac{45-45*1.05^7}{1-1.05}\] \[\Large S_7 = \frac{45-45*1.4071004}{1-1.05}\] \[\Large S_7 = \frac{45-63.319518}{1-1.05}\] \[\Large S_7 = \frac{-18.319518}{-0.05}\] \[\Large S_7 = 366.39036\] The approximate sum of the first 7 terms is 366.39036

OpenStudy (jozelynw):

So the question would be "Find the sum of the first 7 terms

jimthompson5910 (jim_thompson5910):

yes the work shown above would apply to that question

OpenStudy (jozelynw):

Thank you, your name should say lifesaver. You just helped me finish out my math class.

jimthompson5910 (jim_thompson5910):

I'm glad I could be of help